# Wittgenstein’s Family Resemblance Clustering Algorithm

### v1

@@ 1 Introduction
This paper introduces a novel approach in philomatics [^1] Philomatics is a combination of philosophy and mathematics (or machine learning which is a field of mathematics). For more information, refer to (ghojogh2023philomatics). (ghojogh2023philomatics) that draws inspiration from Wittgenstein’s concept of family resemblance in analytic philosophy to develop a clustering algorithm for machine learning. In 1953, in his book Philosophical Investigations, Ludwig Wittgenstein proposed that members of a concept or category are related through overlapping similarities rather than by a single defining property (wittgenstein1953philosophical). According to this view, a family of entities forms a network or chain of interrelated items, where each member resembles some, but not necessarily all, other members. This insight has been widely applied in analytic philosophy to explain the structure of concepts in language, ethics, art, and science, where traditional essentialist definitions fail.
@@
We evaluate the performance of WFR on a variety of toy benchmark datasets. Experimental results demonstrate that the algorithm effectively captures complex, nonlinear cluster structures. These results highlight the potential of WFR as a flexible, philosophically motivated approach to unsupervised learning, bridging insights from analytic philosophy and modern machine learning (ghojogh2023philomatics).
@@ 2.1.1 Biography of Ludwig Wittgenstein
Ludwig Wittgenstein (1889–1951) was an Austrian-British philosopher who made foundational contributions to logic, the philosophy of language, and the philosophy of mind. He studied engineering in Berlin and then moved to Cambridge to study under Bertrand Russell. Later, he became a professor of philosophy at the University of Cambridge. His work profoundly influenced analytic philosophy in the 20th century (wittgenstein1990ludwig).
@@ 2.1.2 Early Wittgenstein (1911–1929)
The early Wittgenstein, active primarily between 1911 and 1929, is exemplified by his work Tractatus Logico-Philosophicus, published in 1921 (wittgenstein1921tractatus). In the Tractatus, he develops a “picture theory” of language, according to which propositions represent the logical structure of reality and determine what can meaningfully be said. The work aims to draw sharp boundaries for language and dissolve philosophical problems by showing them as nonsensical. This book is highly structured, with a hierarchical, numbered system of propositions reminiscent of Spinoza’s Ethics (spinoza1677ethics), and it focuses on the logical form of language and its relation to the world (stanford2002wittgenstein).
@@ 2.1.3 Later Wittgenstein (1930–1951)
After a period away from philosophy, Wittgenstein returned and revised many of his earlier positions. His later philosophy, mainly from 1930 until his death in 1951, is most clearly presented in Philosophical Investigations, published posthumously in 1953 (wittgenstein1953philosophical). In this book, he rejects his primary idea of a single universal logical structure and emphasizes that meaning arises from the use of language in context, introducing the notions of “language-games” and “forms of life.” Philosophy, in the later Wittgenstein, is therapeutic: it clarifies the actual use of language to resolve confusion rather than building a systematic theory (stanford2002wittgenstein).
@@ 2.1.4 Key Differences of Early and Later Wittgenstein
  • Russell’s View: Bertrand Russell explicitly admired Wittgenstein’s early philosophy but strongly criticized his later philosophy (russell1959my).
@@ 2.2 The Notion of Family Resemblance
In Philosophical Investigations (wittgenstein1953philosophical), §§66–71222It is a standard scholarly shorthand, especially in philosophy and law. It means from section 66 through section 71 (inclusive). Note that, in philosophical manuscripts, §  and §§  refer to Section and Sections, respectively., the later Wittgenstein introduced the notion of family resemblance. To illustrate this idea, consider three entities, labeled 1, 2, and 3, as shown in Fig. 1. Suppose entity 1 possesses attributes (or so-called traits) $a$ and $b$, entity 2 possesses attributes $b$ and $c$, and entity 3 possesses attributes $c$ and $d$. Although entities 1 and 3 share no attributes directly, all three may nonetheless be regarded as belonging to the same family. This is because entity 1 shares an attribute with entity 2, and entity 2 shares an attribute with entity 3, forming a chain of overlapping similarities. Thus, resemblance is established not through a single common feature shared by all entities, but through a network or tree-like structure of partial overlaps.
@@ 2.3.1 Ordinary Language Concepts
Family resemblance is invoked when analyzing ordinary language concepts whose extensions are heterogeneous. Wittgenstein’s canonical example is the concept of a game (wittgenstein1953philosophical, §§66–71). Board games, card games, athletic competitions, and solitary amusements share no single common feature. Instead, they exhibit overlapping similarities such as rule-following, competition, skill, or entertainment.
@@ 2.3.2 Philosophy of Language
In analytic philosophy of language, family resemblance is used to support the thesis that meaning is grounded in use rather than in strict definitions (wittgenstein1953philosophical, §43). Words such as “language”, “sentence”, or “meaning” do not admit sharp boundaries but are understood through patterns of use across contexts; thus, they can be understood by family resemblance (wittgenstein1953philosophical, §§65–67).
@@ 2.3.3 Philosophy of Science
Family resemblance has been invoked in discussions of scientific classification, particularly in cases where categories resist essentialist definition (dupre1993disorder; boyd1999homeostasis). A prominent example arises in the classification of biological species. According to essentialist accounts of natural kinds, each genuine kind—such as a biological species or a chemical element—is characterized by an underlying essence, with membership determined by necessary and sufficient conditions (boyd1999homeostasis). However, empirical findings in biology reveal substantial variation among organisms within a species: no single genetic, morphological, or ecological trait is shared by all and only its members, and species boundaries are often indeterminate, as illustrated by phenomena such as ring species, hybridization, and asexual reproduction (mayr1999systematics; dupre1993disorder). In light of this, proponents of a family resemblance approach argue that species are unified not by a common essence but by overlapping clusters of traits, with different members sharing different subsets of these properties (dupre1993disorder; boyd1999homeostasis).
@@ 2.3.4 Aesthetics and Philosophy of Art
In analytic aesthetics, family resemblance is used to reject definitional theories of art. Paintings, musical works, performances, and conceptual art forms share overlapping similarities—such as expression, intention, and cultural role—without a single defining property (weitz2017role; carroll2012philosophy).
@@ 2.3.5 Ethics and Moral Philosophy
Family resemblance has been used in ethics and moral philosophy. Moral concepts such as virtue, responsibility, or vice (opposite of virtue) are often treated as family resemblance concepts. For example, Courage, honesty, generosity, and kindness resemble one another without sharing a common essence (murdoch2013sovereignty). This supports anti-reductive approaches to moral theory (hacker2010virtue).
@@ 2.3.6 Philosophy of Mind
In philosophy of mind and cognitive science, family resemblance is used to analyze mental categories such as emotions or intelligence. Emotions like fear, anger, joy, and shame lack a single shared physiological or intentional structure, instead forming a cluster of related phenomena (ryle2009concept; griffiths2008emotions).
@@ 3.2.3 Kernel Resemblance Function
where $k(.,.)$ is the kernel function and $\phi(.)$ is the pulling function from the input space to the Reproducing Kernel Hilbert Space (RKHS) (ghojogh2023background).
@@ 3.3.1 Nearest Neighbors Graph
We first find the $k$-Nearest Neighbors (kkNN) of the training data instances. Thus, the neighbors of each data instance is found and a kkNN graph is formed. Different algorithms can be used for finding the kkNN graph (bhatia2010survey). Some of the possible algorithms to use are brute-force (exhaustive) nearest neighbor, KD-Tree nearest neighbor (friedman1977algorithm), and Ball Tree nearest neighbor (omohundro1989five; dolatshah2015ball), which are implemented in Scikit-learn library (pedregosa2011scikit).
@@ 3.3.3 Search in the Adjacency Graph
The adjacency matrix $\widetilde{\boldsymbol{A}}$ induces an adjacency graph in which adjacent data instances are connected. Graph search algorithms (cormen2022introduction), such as Depth-First Search (DFS) (tarjan1972depth) or Breadth-First Search (BFS) (moore1959shortest), can be employed to identify the connected components of this graph. Each connected component corresponds to a family, or cluster, in which data instances are related through a chain of resemblance. Binary connectivity captures the existence of family resemblance chains rather than their strength.
@@ 3.5 Automatic Thresholding
Clustering is inherently an ill-defined problem, as the perceived number of clusters can vary depending on the observer’s perspective. For instance, one person may identify two clusters in a dataset, while another may perceive three clusters by examining finer separations. Consequently, all clustering algorithms involve at least one hyperparameter that determines or influences the number of clusters. Some algorithms, such as K-means (mcqueen1967some), explicitly require the number of clusters as input, whereas others, such as DBSCAN (ester1996density), include hyperparameters that indirectly affect the number of clusters.
@@
Note that it is also possible to use alternative clustering evaluation scores, such as Silhouette score (rousseeuw1987silhouettes) or Davies–Bouldin index (davies2009cluster).
@@ 4.2 The Complexity of Test Phase
K-means (mcqueen1967some) $\mathcal{O}(ncdI)$ $\mathcal{O}(nd)$
@@
Affinity Propagation (frey2007clustering) $\mathcal{O}(n^{2}I)$ $\mathcal{O}(n^{2})$
Mean Shift (cheng1995mean) $\mathcal{O}(n^{2}I)$ $\mathcal{O}(n)$
Spectral Clustering (ng2001spectral) $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$
@@
Ward Clustering (ward1963hierarchical) $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$
DBSCAN (ester1996density) $\mathcal{O}(n\log n)$ $\mathcal{O}(n)$
OPTICS (ankerst1999optics) $\mathcal{O}(n\log n)$ $\mathcal{O}(n)$
HDBSCAN (campello2015hierarchical) $\mathcal{O}(n\log n)$ $\mathcal{O}(n)$
BIRCH (zhang1996birch) $\mathcal{O}(n)$ $\mathcal{O}(n)$
@@ 4.3 Comparison with Baseline Clustering Algorithms
Table 1 compares the computational complexity of WFR clustering with several well-known clustering algorithms. As shown in this table, many classical clustering algorithms incur quadratic or cubic time and space complexity, which limits their scalability to large datasets. Spectral clustering (ng2001spectral; shi2000normalized; ghojogh2023laplacian), affinity propagation (frey2007clustering), and hierarchical methods (murtagh2012algorithms; johnson1967hierarchical) require $\mathcal{O}(n^{2})$ memory, while spectral clustering additionally requires $\mathcal{O}(n^{3})$ time due to eigen-decomposition.
Density-based methods such as DBSCAN (ester1996density), OPTICS (ankerst1999optics), and HDBSCAN (campello2015hierarchical) achieve near-linear time complexity under suitable indexing assumptions, but rely on density estimation and distance thresholds that may be sensitive to data distribution and scaling. Centroid-based methods such as K-means (mcqueen1967some) and Gaussian mixture models (mclachlan2000finite; dempster1977maximum; ghojogh2019fitting) scale well but require the number of clusters to be specified in advance and struggle with nonconvex cluster shapes.
@@ 5.2 Comparison with Other Clustering Algorithms
In this section, we compare our proposed WFR clustering algorithm with well-known, effective clustering algorithms, including K-means (mcqueen1967some), affinity propagation (frey2007clustering), mean shift (cheng1995mean; comaniciu2002mean), spectral clustering (ng2001spectral; shi2000normalized; ghojogh2023laplacian), Ward (ward1963hierarchical), agglomerative clustering (hierarchical clustering) (murtagh2012algorithms; johnson1967hierarchical), DBSCAN (ester1996density), HDBSCAN (campello2015hierarchical), OPTICS (ankerst1999optics), BIRCH (zhang1996birch), and gaussian mixture models using Expectation Maximization (EM) (mclachlan2000finite; dempster1977maximum; ghojogh2019fitting). For the baseline clustering methods, we use their optimized libraries available in Scikit-learn (pedregosa2011scikit). A Scikit-learn–compatible implementation of the proposed WFR algorithm is available in the following GitHub repository: https://github.com/bghojogh/WFR_Clustering.

### v2

@@ 1 Introduction
This paper introduces a novel approach in philomatics [^1] Philomatics is a combination of philosophy and mathematics (or machine learning which is a field of mathematics). For more information, refer to (Ghojogh and Babaie, 2023). (Ghojogh and Babaie, 2023) that draws inspiration from Wittgenstein’s concept of family resemblance in analytic philosophy to develop a clustering algorithm for machine learning. In 1953, in his book Philosophical Investigations, Ludwig Wittgenstein proposed that members of a concept or category are related through overlapping similarities rather than by a single defining property (Wittgenstein, 1953). According to this view, a family of entities forms a network or chain of interrelated items, where each member resembles some, but not necessarily all, other members. This insight has been widely applied in analytic philosophy to explain the structure of concepts in language, ethics, art, and science, where traditional essentialist definitions fail.
@@
We evaluate the performance of WFR on a variety of toy benchmark datasets. Experimental results demonstrate that the algorithm effectively captures complex, nonlinear cluster structures. These results highlight the potential of WFR as a flexible, philosophically motivated approach to unsupervised learning, bridging insights from analytic philosophy and modern machine learning (Ghojogh and Babaie, 2023).
@@ 2.1.1 Biography of Ludwig Wittgenstein
Ludwig Wittgenstein (1889–1951) was an Austrian-British philosopher who made foundational contributions to logic, the philosophy of language, and the philosophy of mind. He studied engineering in Berlin and then moved to Cambridge to study under Bertrand Russell. Later, he became a professor of philosophy at the University of Cambridge. His work profoundly influenced analytic philosophy in the 20th century (Monk, 1990).
@@ 2.1.2 Early Wittgenstein (1911–1929)
The early Wittgenstein, active primarily between 1911 and 1929, is exemplified by his work Tractatus Logico-Philosophicus, published in 1921 (Wittgenstein, 1921). In the Tractatus, he develops a “picture theory” of language, according to which propositions represent the logical structure of reality and determine what can meaningfully be said. The work aims to draw sharp boundaries for language and dissolve philosophical problems by showing them as nonsensical. This book is highly structured, with a hierarchical, numbered system of propositions reminiscent of Spinoza’s Ethics (Spinoza, 1677), and it focuses on the logical form of language and its relation to the world (Stanford Encyclopedia of Philosophy, 2002).
@@ 2.1.3 Later Wittgenstein (1930–1951)
After a period away from philosophy, Wittgenstein returned and revised many of his earlier positions. His later philosophy, mainly from 1930 until his death in 1951, is most clearly presented in Philosophical Investigations, published posthumously in 1953 (Wittgenstein, 1953). In this book, he rejects his primary idea of a single universal logical structure and emphasizes that meaning arises from the use of language in context, introducing the notions of “language-games” and “forms of life.” Philosophy, in the later Wittgenstein, is therapeutic: it clarifies the actual use of language to resolve confusion rather than building a systematic theory (Stanford Encyclopedia of Philosophy, 2002).
@@ 2.1.4 Key Differences of Early and Later Wittgenstein
  • Russell’s View: Bertrand Russell explicitly admired Wittgenstein’s early philosophy but strongly criticized his later philosophy (Russell, 1959).
@@ 2.2 The Notion of Family Resemblance
In Philosophical Investigations (Wittgenstein, 1953), §§66–71222It is a standard scholarly shorthand, especially in philosophy and law. It means from section 66 through section 71 (inclusive). Note that, in philosophical manuscripts, §  and §§  refer to Section and Sections, respectively., the later Wittgenstein introduced the notion of family resemblance. To illustrate this idea, consider three entities, labeled 1, 2, and 3, as shown in Fig. 1. Suppose entity 1 possesses attributes (or so-called traits) $a$ and $b$, entity 2 possesses attributes $b$ and $c$, and entity 3 possesses attributes $c$ and $d$. Although entities 1 and 3 share no attributes directly, all three may nonetheless be regarded as belonging to the same family. This is because entity 1 shares an attribute with entity 2, and entity 2 shares an attribute with entity 3, forming a chain of overlapping similarities. Thus, resemblance is established not through a single common feature shared by all entities, but through a network or tree-like structure of partial overlaps.
@@ 2.3.1 Ordinary Language Concepts
Family resemblance is invoked when analyzing ordinary language concepts whose extensions are heterogeneous. Wittgenstein’s canonical example is the concept of a game (Wittgenstein, 1953, §§66–71). Board games, card games, athletic competitions, and solitary amusements share no single common feature. Instead, they exhibit overlapping similarities such as rule-following, competition, skill, or entertainment.
@@ 2.3.2 Philosophy of Language
In analytic philosophy of language, family resemblance is used to support the thesis that meaning is grounded in use rather than in strict definitions (Wittgenstein, 1953, §43). Words such as “language”, “sentence”, or “meaning” do not admit sharp boundaries but are understood through patterns of use across contexts; thus, they can be understood by family resemblance (Wittgenstein, 1953, §§65–67).
@@ 2.3.3 Philosophy of Science
Family resemblance has been invoked in discussions of scientific classification, particularly in cases where categories resist essentialist definition (Dupré, 1993; Boyd and others, 1999). A prominent example arises in the classification of biological species. According to essentialist accounts of natural kinds, each genuine kind—such as a biological species or a chemical element—is characterized by an underlying essence, with membership determined by necessary and sufficient conditions (Boyd and others, 1999). However, empirical findings in biology reveal substantial variation among organisms within a species: no single genetic, morphological, or ecological trait is shared by all and only its members, and species boundaries are often indeterminate, as illustrated by phenomena such as ring species, hybridization, and asexual reproduction (Mayr, 1999; Dupré, 1993). In light of this, proponents of a family resemblance approach argue that species are unified not by a common essence but by overlapping clusters of traits, with different members sharing different subsets of these properties (Dupré, 1993; Boyd and others, 1999).
@@ 2.3.4 Aesthetics and Philosophy of Art
In analytic aesthetics, family resemblance is used to reject definitional theories of art. Paintings, musical works, performances, and conceptual art forms share overlapping similarities—such as expression, intention, and cultural role—without a single defining property (Weitz, 2017; Carroll, 2012).
@@ 2.3.5 Ethics and Moral Philosophy
Family resemblance has been used in ethics and moral philosophy. Moral concepts such as virtue, responsibility, or vice (opposite of virtue) are often treated as family resemblance concepts. For example, Courage, honesty, generosity, and kindness resemble one another without sharing a common essence (Murdoch and Midgley, 2013). This supports anti-reductive approaches to moral theory (Hacker-Wright, 2010).
@@ 2.3.6 Philosophy of Mind
In philosophy of mind and cognitive science, family resemblance is used to analyze mental categories such as emotions or intelligence. Emotions like fear, anger, joy, and shame lack a single shared physiological or intentional structure, instead forming a cluster of related phenomena (Ryle and Tanney, 2009; Griffiths, 2008).
@@ 3.2.3 Kernel Resemblance Function
where $k(.,.)$ is the kernel function and $\phi(.)$ is the pulling function from the input space to the Reproducing Kernel Hilbert Space (RKHS) (Ghojogh et al., 2023a).
@@ 3.3.1 Nearest Neighbors Graph
We first find the $k$-Nearest Neighbors (kkNN) of the training data instances. Thus, the neighbors of each data instance is found and a kkNN graph is formed. Different algorithms can be used for finding the kkNN graph (Bhatia and Vandana, 2010). Some of the possible algorithms to use are brute-force (exhaustive) nearest neighbor, KD-Tree nearest neighbor (Friedman et al., 1977), and Ball Tree nearest neighbor (Omohundro, 1989; Dolatshah et al., 2015), which are implemented in Scikit-learn library (Pedregosa et al., 2011).
@@ 3.3.3 Search in the Adjacency Graph
The adjacency matrix $\widetilde{\boldsymbol{A}}$ induces an adjacency graph in which adjacent data instances are connected. Graph search algorithms (Cormen et al., 2022), such as Depth-First Search (DFS) (Tarjan, 1972) or Breadth-First Search (BFS) (Moore, 1959), can be employed to identify the connected components of this graph. Each connected component corresponds to a family, or cluster, in which data instances are related through a chain of resemblance. Binary connectivity captures the existence of family resemblance chains rather than their strength.
@@ 3.5 Automatic Thresholding
Clustering is inherently an ill-defined problem, as the perceived number of clusters can vary depending on the observer’s perspective. For instance, one person may identify two clusters in a dataset, while another may perceive three clusters by examining finer separations. Consequently, all clustering algorithms involve at least one hyperparameter that determines or influences the number of clusters. Some algorithms, such as K-means (McQueen, 1967), explicitly require the number of clusters as input, whereas others, such as DBSCAN (Ester et al., 1996), include hyperparameters that indirectly affect the number of clusters.
@@
Note that it is also possible to use alternative clustering evaluation scores, such as Silhouette score (Rousseeuw, 1987) or Davies–Bouldin index (Davies and Bouldin, 2009).
@@ 4.2 The Complexity of Test Phase
K-means (McQueen, 1967) $\mathcal{O}(ncdI)$ $\mathcal{O}(nd)$
@@
Affinity Propagation (Frey and Dueck, 2007) $\mathcal{O}(n^{2}I)$ $\mathcal{O}(n^{2})$
Mean Shift (Cheng, 1995) $\mathcal{O}(n^{2}I)$ $\mathcal{O}(n)$
Spectral Clustering (Ng et al., 2001) $\mathcal{O}(n^{3})$ $\mathcal{O}(n^{2})$
@@
Ward Clustering (Ward Jr, 1963) $\mathcal{O}(n^{2})$ $\mathcal{O}(n^{2})$
DBSCAN (Ester et al., 1996) $\mathcal{O}(n\log n)$ $\mathcal{O}(n)$
OPTICS (Ankerst et al., 1999) $\mathcal{O}(n\log n)$ $\mathcal{O}(n)$
HDBSCAN (Campello et al., 2015) $\mathcal{O}(n\log n)$ $\mathcal{O}(n)$
BIRCH (Zhang et al., 1996) $\mathcal{O}(n)$ $\mathcal{O}(n)$
@@ 4.3 Comparison with Baseline Clustering Algorithms
Table 1 compares the computational complexity of WFR clustering with several well-known clustering algorithms. As shown in this table, many classical clustering algorithms incur quadratic or cubic time and space complexity, which limits their scalability to large datasets. Spectral clustering (Ng et al., 2001; Shi and Malik, 2000; Ghojogh et al., 2023b), affinity propagation (Frey and Dueck, 2007), and hierarchical methods (Murtagh and Contreras, 2012; Johnson, 1967) require $\mathcal{O}(n^{2})$ memory, while spectral clustering additionally requires $\mathcal{O}(n^{3})$ time due to eigen-decomposition.
Density-based methods such as DBSCAN (Ester et al., 1996), OPTICS (Ankerst et al., 1999), and HDBSCAN (Campello et al., 2015) achieve near-linear time complexity under suitable indexing assumptions, but rely on density estimation and distance thresholds that may be sensitive to data distribution and scaling. Centroid-based methods such as K-means (McQueen, 1967) and Gaussian mixture models (McLachlan and Peel, 2000; Dempster et al., 1977; Ghojogh et al., 2019) scale well but require the number of clusters to be specified in advance and struggle with nonconvex cluster shapes.
@@ 5.2 Comparison with Other Clustering Algorithms
In this section, we compare our proposed WFR clustering algorithm with well-known, effective clustering algorithms, including K-means (McQueen, 1967), affinity propagation (Frey and Dueck, 2007), mean shift (Cheng, 1995; Comaniciu and Meer, 2002), spectral clustering (Ng et al., 2001; Shi and Malik, 2000; Ghojogh et al., 2023b), Ward (Ward Jr, 1963), agglomerative clustering (hierarchical clustering) (Murtagh and Contreras, 2012; Johnson, 1967), DBSCAN (Ester et al., 1996), HDBSCAN (Campello et al., 2015), OPTICS (Ankerst et al., 1999), BIRCH (Zhang et al., 1996), and gaussian mixture models using Expectation Maximization (EM) (McLachlan and Peel, 2000; Dempster et al., 1977; Ghojogh et al., 2019). For the baseline clustering methods, we use their optimized libraries available in Scikit-learn (Pedregosa et al., 2011). A Scikit-learn–compatible implementation of the proposed WFR algorithm is available in the following GitHub repository: https://github.com/bghojogh/WFR_Clustering.
@@ 6 Conclusion and Future Directions
Acknowledgement
The authors were unaware of the related work in (Xiao and Yu, 2010) during the development of the WFR algorithm and the preparation of the manuscript. While related in spirit, their method differs from WFR clustering in several aspects. Interested readers may consult (Xiao and Yu, 2010) for an alternative family-resemblance-based clustering approach. Connections between family resemblance and machine learning have also been discussed in informal venues (Philosophy Stackexchange, 2017). Family resemblance has further been explored in the context of fuzzy sets (Veri, 2023).